OpenStudy (anonymous):

My answer is C Which of the following is a polynomial with roots -4, -2 i , and 2 i ? (1 points) Answer x3 - 4x2 + 4x - 16 x3 + 4x2 + 4x + 16 x3 + 2x2 + 4x + 8 x3 - 2x2 + 4x - 8

OpenStudy (amoodarya):

b is correct put 4 in them then put 2i you will find

OpenStudy (anonymous):

oh let me try:)

OpenStudy (whpalmer4):

If we have those roots, the factors of the polynomial will be \[(x+4)(x+2i)(x-2i)\] because the roots are the points where the polynomial equals 0 which means that one or more of the factors equals 0. If multiplied out, we get: \[(x+4)(x+2i)(x-2i) = (x+4)(x^2-2ix+2ix-4i^2) = \]\[(x+4)(x^2+4) = x^3+4x+4x^2+16 = x^3+4x^2+4x+16\]

OpenStudy (anonymous):

i got b!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (whpalmer4):

In real life, you may not get to choose from a list of answers, so it's good to be able to do the problem for real, not just game the test by eliminating wrong answers :-)

OpenStudy (whpalmer4):

Though I agree that ability to game the test is a very useful skill, too!

OpenStudy (anonymous):

you're right, I appreciate your input